# This Candidate is [MASK]

### v1

@@ 4.2 Nonlinearities
Note: Partial dependence contours depict average job market outcomes when the whole sample is assigned (counterfactual) levels of average letter polarity and length.
A final point involves interaction between letter sentiment and letter length. Figure 10 shows that, predictably, these two forms of signalling positively interact
5 Preliminary Conclusions
There is compelling evidence that reference letters are predictive of job market outcomes. Both average letter length and sentiment polarity are positively associated with the probability of success in general research positions. The relationship seems to be even stronger for more selective “elite” positions.
All in all, the evidence seems to be consistent with the canonical Saloner (1985) model. There is competition among referees to place their students in good places, and this creates incentives to somewhat “inflate” candidates’ achievements and prospects. However, the repetitive nature of the interaction between letter-writers and employers, plus the fact that individual ability is eventually revealed mean that there are counterbalancing forces. The resulting equilibrium is one in which relatively stronger candidates get better recommendations, employers place significant trust in the contents of the references, and those candidates are more likely to get the best job offers.
This meritocratic mechanism seems to operate quite broadly, across fields of research, personal characteristics of the candidate and throughout the rankings. In fact, one of the most refreshing takeaways from the findings is that good candidates from relatively lower ranked PhDs can offset a sizeable part of any starting point disadvantage by producing research worthy of excellent referee reports.
@@ 5 Preliminary Conclusions
On the methodological front, a key takeaway of the evidence presented here is that dictionary methods have limited power. Because they do not capture context, they are unable to extract some relevant information. The prompt-based strategy described here is a simple enough alternative that leverages the enormous resources invested in pre-training LLMs. Extending and refining this method should be the focus of further research.
@@ Appendix A Appendix
Table A.1: Sub-group Sentiment Polarity and Letter Length
BB Prompt-based Sentiment (pp.) Word Count (,000s)
Period
2013 0.262 0.845
2015 0.250 0.949
2016 0.236 0.880
2017 0.229 0.977
2018 0.253 0.924
2019 0.259 1.060
2020 0.265 1.058
2021 0.280 1.197
p-val 0.164 0.000
Gender
Female 0.277 0.982
Male 0.249 1.003
p-val 0.006 0.458
Field
Applied 0.261 1.042
Econometrics 0.235 0.959
Finance 0.264 0.896
Macroeconomics 0.279 1.000
Theory 0.174 1.144
p-val 0.000 0.000
PhD Rank
1–25 0.238 1.077
26–50 0.255 1.036
51–75 0.257 0.937
76–100 0.255 0.972
101–150 0.253 0.988
151–200 0.277 1.003
201+ 0.283 0.836
p-val 0.109 0.000
Citizenship Region
Asia 0.247 0.948
Developing 0.238 0.998
Europe 0.270 1.044
Former Soviet 0.248 0.995
North America 0.296 1.007
p-val 0.056 0.045
PhD Region
USA 0.253 0.989
English Speaking 0.275 0.991
Europe 0.252 1.021
Other 0.252 0.864
p-val 0.296 0.376
Notes: The p-val correspond to an anova test of equality of means across groups.
@@
Permanent Research Position Elite Position
Max Depth 4 5
Min Samples to Split Node 30 29
Min Samples in Leaf Node 13 9
Forest Size 140 150
Notes: Hyper-parameters result from 5-fold CV search. Samples are bootstrapped and $\sqrt{9}$ predictors are selected at random in the CART tree estimation. The impurity measure is the Gini.

### v2

@@ 4.3 Random Forest
Notes: Partial dependence contours simulate average job market outcomes when the whole sample is assigned counterfactual levels of two continuous predictors simultaneously. The level curves depict different probabilities of success in the academic job market.
Figure 9: Partial Dependence Interactions
(a) Sex
(b) Candidate Region
(c) Field of Research
(d) PhD Rank
Note: Partial dependence interactions simulate average job market outcomes when the whole sample is assigned counterfactual values for Avg. RL sentiment and one discrete predictor.
The plots in figure 9 present similar simulations, this time by assigning counterfactual levels to Avg. RL sentiment and four different discrete predictors: [^43] Similar plots for interactions with Avg. RL length and Sentiment Range are omitted to save space. These and other $\mathrm{pd}$ plots are available in the online appendix (TODO).
  • (a)
    Female candidates in my sample generally perform better than males. Everyone benefits from positive RLs at more or less the same rate. In section 5 below, I show that the real picture is less rosy when considering other aspects of RL content.
  • (b)
    There is a clear pattern that seems to strongly disfavor candidates from Asia. Not only do they have a lower probability of success, but also seem to benefit less from positive RLs.
  • (c)
    There seems to be a small difference in rates of success across fields of research for low average sentiment levels. Finance candidates perform a bit better than applied people and theorists. However, these differences become smaller as RL sentiment increases.
  • (d)
    In contrast, the random forest predictions suggest that the ranking of the PhD program makes a very sizeable difference in terms of obtaining a research position. Specifically, candidates from the top-25 programs have markedly higher success rates, and the gap is strongly increasing with RL sentiment. Candidates from these programs that manage to secure RLs in the top decile of sentiment have an almost 60% predicted probability of attaining a position in a top research institution.
4.4 Comparison of Prompt-based Sentiment with Other Approaches
Prompt-based sentiment applied to RLs can predict success in the job market. Can other approaches reproduce this feat? To answer this question I obtained measures of sentiment polarity using each of the approaches discussed in section 2.444444The HGI and LM dictionaries can be applied directly to RLs. Vader is a sentence-based sentiment analyzer, so I apply it sentence-by-sentence to every RL and then average the resulting scores. Flair and FinBert are language models with relatively small context windows (512 tokens). In order to obtain sentiment scores for the RLs, I break them into chunks and then average the scores. Table A.1 in the appendix presents descriptive statistics for the different sentiment measures.
Figure 10: Correlation Matrix
Notes: NN=1968. Pearson correlation coefficients among all RL sentiment measures.
Figure 10 presents Pearson correlations among the different sentiment measures. While almost all correlations are positive, it is noteworthy that the coefficients are clustered by technique. For example, lexicon-based measures have much stronger associations among each other—correlations close to 0.5—than with any of the LLM-based measures. Similarly, the correlation between the two instruction-based measures is 0.82.
I repeated the regression analysis from section 4.1 for each of the alternative sentiment measures. Figure 11 summarizes my findings. Specifically, the box plots in the figure show the p-value distributions for the Avg. RL Sentiment coefficient in equation (3). For each sentiment measure, I estimated 150 different specifications by varying the outcome variable (5 variants), considering alternative control variable sets (5), using either the full or the restricted ‘complete applications sample’ (2), and applying different assumptions to the error term (3).
Somewhat predictably, neither the lexical approach nor the fine-tuned LLMs are able to predict job market outcomes. The dictionaries involved in the former and the fine-tuning process in the latter target applications that are quite distant from RL sentiment extraction.
As mentioned in section 2, ModernBert is a bidirectional LLM (encoder). These models are smaller and relatively light-weight compared to the generative LLMs. They also allow for a more flexible prompt strategy given that the [MASK] token can be included anywhere in the template. However, they suffer from a limited context window. In the case of ModernBert, the 8K token context is just enough to fit every RL in one go. However, using the full context window comes at a cost in performance. Figure A.2 (left) shows the distribution of sentiment scores when applying the prompt-based strategy using ModernBert. Comparing this figure to figure 1, we can see that ModernBert produces scores that are decreasing in the length of the RL, which is the unfortunate consequence of the context size limitation.
Regardless of these limitations, the sentiment scores that result from applying ModernBert can contribute to predicting job market outcomes. In most of the regressions, the average RL sentiment comes out significant at the 5% level. [^45] I also trained a random forest using the ModernBert score and found that it was the most important feature in predicting the baseline outcome.
Figure 11: Predicting Job Market Outcomes — Alternative Sentiment Approaches
Notes: The box plots depict the p-value distributions for the coefficient corresponding to Avg. RL Sentiment in equation (3). Each sentiment measure is used in 150 alternative specifications: 5 outcome variables, 2 samples, 5 control variable groups, and 3 alternative clusterings of the errors.
The two instruction-based sentiment scores are also successful in predicting outcomes, with Llama 4 clearly outperforming DeepSeek 3 in this exercise. In figure A.2, I also plot the distributions of these sentiment measures against the RL word count. In this case, the intuitive positive relationship between RL length and sentiment reappears. However, the instruction-based measures also have disadvantages. It is clear from the figure that the model generates answers that are “round” numbers. In addition, creating a prompt that produced sufficient variation among candidates took many repeated trials. In particular, as is clear from the instruction set I settled for, it was quite difficult to get the model to generate low scores for at least some RLs.
Figure 12: The Role of Sentiment Range — Alternative Sentiment Approaches
Notes: The box plots depict the p-value distributions for the coefficients corresponding to Avg. RL Sentiment and RL Sentiment Range in equation (4). Each sentiment measure is used in 75 alternative specifications estimated on the ‘full applications’ restricted sample: 5 outcome variables, 5 control variable groups, and 3 alternative clusterings of the errors.
Figure 12 repeats the exercise, this time for equation (4). All regressions are estimated on the ‘full application’ restricted sample. Neither ModernBert nor the instruction-based models are able to replicate the findings in table 10.
5 Gendered Sentiment
Prompt-based learning is a general technique for text analysis. In this section, I show how it can be deployed to extract information regarding the candidate’s main strengths, as evaluated by the letter writer.
There is a large literature specialized on seemingly harmless differences in how women and men are described. Within economics, Eberhardt et al. (2023), Baltrunaite et al. (2024), and Hochleitner et al. (2025) have shown that the language used to describe women in RLs accentuates what they call ‘grindstone’ aspects of their personality. These are traits such as conscientiousness, patience, attention to detail, being a good communicator and collaborator, etc. Male candidates, on the other hand, are more often described with ‘standout’ traits such as originality, leadership, technical excellence, etc. Theses small differences in language are not really harmless in their effects in the job market.
The papers in this literature mostly follow a lexical approach. [^46] The approach in Baltrunaite et al. (2024) is somewhat different in that they use embeddings. They have curated dictionaries of terms associated with personality traits, and then classified RLs based on the relative frequency with which terms in the different dictionary categories are used in the RL text. In this section, I show how a prompt-based strategy can be used to extract analogous information.
Table 12: Prompt-based Extraction of Personality Traits
Prompt ($\mathcal{T}_{personality}$) Verbalizer ($\mathcal{V}_{personality}$)
“[X] If I had to pick this candidate’s main strength, I would mention their [MASK]” standout: {ability, research, creativity, originality, leadership, skill, skills, intelligence, theoretical, technical, mathematical, analytical, statistical, intellectual, original, critical, ambition, drive, excellence} grindstone: {passion, dedication, diligence, patience, perseverance, persistence, open, personality, genuine, interpersonal, collaborative, enthusiasm, work, attention, humility, communication, maturity, flexibility, commitment, writing, teaching}
Table 12 presents the template/verbalizer combination that I employed. A key distinctive feature of my approach is that the model is implicitly forced to choose which aspect of the candidate’s personality to emphasize. For the sake of brevity, throughout this section I focus mostly on the difference between the probabilities assigned to the two types of traits:
$\mathrm{P}\left(y=\text{standout}\mid\boldsymbol{x}\right)-\mathrm{P}\left(y=\text{grindstone}\mid\boldsymbol{x}\right).$
For lack of a better term, I refer to this quantity as the net-standout score.
Figure A.3 in the appendix shows the distribution of the net-standout score. The distribution is symmetrical and has a remarkably wide range, going from strongly negative (grindstone-biased) to highly positive (strongly standout-biased) RLs. [^47] Summary statistics for the net-standout score can be found in table A.1 in the appendix. Similarly to the prompt-based sentiment polarity measure, net-standout is clearly increasing in RL length. Remarkably, however, the scatter plot of sentiment polarity and net-standout has no distinguishable pattern. The correlation coefficient between these two scores is 0.13. In summary, net-standout measures a completely independent feature of the information contained in RLs.
Figure 13: Net-Standout Score — Gender Differences
Notes: NN=1968. The box plots show the distribution of the net-standout score for female and male candidates.
Are there differences in net-standout scores between male and female candidates? As is clear from figure 13, the answer is a rotund ‘yes’. Similar to the literature, I find that RLs written for male candidates emphasize ‘standout’ traits more, whereas females are more likely to be described using ‘grindstone’ attributes. Moreover, the gender effect is robust to introducing controls for other candidates characteristics, for the PhD rank and region, for the job market year, and for a host of letter writer characteristics including sex, whether the writer is the main adviser, whether the writer is either a top-10 or a top-5 percentile author, and the writer’s title and potential experience. [^48] These regression results are available in the online appendix (TODO).
Figure 14: Predicting Net-Standout — Partial Dependence Tables
Note: Partial dependence tables simulate the RL net-standout score when the whole sample is assigned counterfactual values for the candidate’s sex and one additional discrete predictor.
Figure 15: Predicting Net-Standout — Letter Writer Experience
Note: Partial dependence interaction simulates the RL net-standout score when the whole sample is assigned counterfactual values for the candidate’s sex and the letter writer’s potential experience (in decades).
Using a random forest, I once again explore interactions, this time between the candidate’s sex and other potential predictors of the net-standout score. [^49] Full details on this procedure, including hyper-parameters and performance metrics are available in the online appendix (TODO). The top panel of figure 14 shows the interaction between candidate sex and the rank of their PhD program. The gender gap in net-standout is present throughout the ranking distribution. The table implies that a male candidate from a PhD program ranked 76–100 is likely to get a higher net-standout score than a female in a top-25 program.
The table in the bottom right panel reinforces this point. Regardless of the prestige of the letter writer, women are more likely to be described with grindstone terms. Finally, the bottom left panel shows that female candidates cannot expect to receive a higher net-standout score if they pick a female referee.
Figure 15 is meant to provide some hint as to whether we can expect this gender gap to disappear in the near future. It shows the interaction of candidate sex and letter writer potential experience (in decades) in explaining the net-standout score. While younger scholars seem to be more inclined to emphasize candidates’ ability over other traits, there is no evidence that the gap is any narrower for letter writers who received their PhDs recently.
In appendix table A.6, I present regression results when equation (3) is augmented to include ‘gendered’ sentiment variables. Specifically, I include the average standout ($\mathrm{P}\left(y=\text{standout}\mid\boldsymbol{x}\right)$) and average grindstone ($\mathrm{P}\left(y=\text{grindstone}\mid\boldsymbol{x}\right)$) scores as predictors of job market outcomes. The coefficients for average letter length and sentiment are almost identical to those in table 8. The coefficients for the average standout score are relatively small and statistically insignificant. The average grindstone score, on the other hand, has a negative and statistically significant association with the probability of success. [^50] The exception is again the outcome that excludes post-docs from the definition of success.
6 Conclusion
@@
The lexical approach to sentiment analysis has limited power. Because dictionaries cannot capture context, the resulting sentiment measures are unable to extract some relevant information. The prompt-based strategy described in this paper is a simple alternative that leverages the enormous resources invested in pre-training LLMs. Prompt-based sentiment extraction offers numerous advantages:
  1. 1.
    The text input can be entered “as is”, without the need of any heavy-handed pre-processing. Multiple languages are acceptable.
  2. 2.
    LLMs take context into account.
  3. 3.
    No labeled data and no fine-tuning is necessary.
  4. 4.
    The resulting sentiment scores have a probability interpretation.
  5. 5.
    The model performance can be benchmarked against known meta-data.
  6. 6.
    The technique can be extended to capture text features other than general sentiment.
In addition to this “static” advantages, it should be kept in mind that LLMs are becoming more powerful at a very fast pace. As a result, the advantages of prompt-based learning over the alternatives can only be expected to grow over time.
However, this approach to sentiment analysis is not without its challenges. Prompt and verbalizer engineering can be time-consuming. For certain applications, coming up with an appropriate prompt may be challenging. And implementing prompt-based learning is certainly not a trivial programming task.
Some researchers will be more inclined to simply query an instruction-based model à la Chat-GPT. I show that the output from such an exercise has to always be taken with a pinch of salt. These models pass the full output vector to a text-generating function that effectively samples a single token from the whole vocabulary.
In order to showcase the technique, in this paper I apply prompt-based sentiment extraction to a small administrative dataset of confidential RLs. As a form of validation, I show that relatively simple prompts can predict known candidate meta-data such as sex and field of research.
Using the resulting sentiment scores, I show that there is compelling evidence that RLs are predictive of job market outcomes. The candidate’s average letter polarity is strongly predictive of success in the job market, regardless of the precise definition of the outcome variable. Moreover, I show that disagreement among the letter writers is penalized by the market.
All in all, the evidence seems to be roughly consistent with the canonical Saloner (1985) model. While there is competition among referees to place their students in good institutions, the repetitive nature of the interaction between letter-writers and employers means that there are counterbalancing forces. The resulting equilibrium is one in which relatively stronger candidates get better recommendations, employers place significant trust in the contents of the RLs, and those candidates are more likely to get the best job offers. This meritocratic mechanism seems to operate quite broadly, across fields of research, personal characteristics of the candidate and throughout the rankings (both of PhD programs and of letter writers).
On a less positive note, using a different prompt I am able to show that there is a persistent gap in how male and female candidates are described in RLs. Specifically, RLs written for females tend to be somewhat biased towards ‘grindstone’ personality traits. This gender gap, which exists regardless of field, PhD program rank and letter-writer characteristics, is likely one of the contributing factors to the continuing under-representation of women in the profession.
@@ Appendix A Appendix
Table A.1: Descriptive Statistics — Alternative Sentiment Measures
Lexical Approach Fine-tuned LLMs
Harvard Loughran Vader Flair FinBert
G.I. McDonald
Mean 0.516 0.088 0.257 0.982 0.396
Std. Dev. 0.157 0.304 0.098 0.098 0.445
Min -0.089 -1.000 -0.153 0.004 -0.826
Median 0.520 0.080 0.259 1.000 0.094
Max 1.000 1.000 0.740 1.000 1.000
Prompt-based Sentiment Instruction-tuned LLMs
Llama Modern Net-Standout Llama Instruct DeepSeek
3.1 Bert Score 4/16E 3
Mean 0.073 0.282 0.176 0.696 0.822
Std. Dev. 0.034 0.167 0.076 0.219 0.094
Min -0.008 0.000 -0.145 0.000 0.200
Median 0.068 0.256 0.179 0.800 0.800
Max 0.221 0.868 0.441 1.000 0.900
Notes: NN=1968. See Section 2 for an explanation of each of these approaches to sentiment extraction. The net-standout score is introduced in section 5.
Table A.2: Predicting Success in the Academic Job Market: control variables
Baseline Outcome: top-100 academic and top-50 policy jobs
(model 5)
Sex: Male -0.1155∗∗∗ PhD Region: USA -0.0623
(0.0393) [0.0401] ||0.0444|| (0.0367) [0.0320] ||0.0361||
Candidate’s Region (ref. Asia) Field (ref. Applied)
Developing Countries 0.0953 Econometrics 0.1411
[…]
Europe 0.0507 Finance 0.1694∗∗∗
(0.0415) [0.0377] ||0.0377|| (0.0447) [0.0450] ||0.0448||
Former USSR 0.0653 Macroeconomics 0.1050
[…]
North America 0.2204∗∗ Theory -0.0725
(0.1035) [0.1035] ||0.0952|| (0.0521) [0.0560] ||0.0585||
Publication(s) 0.0964 PhD Length 0.0192
(0.0725) [0.0720] ||0.0648|| $\leq$6 years (0.0553) [0.0559] ||0.0436||
PhD University Rank (ref. 1–25) Top-5% Letter Writers (ref. 0)
26–50 -0.1345 One -0.0017
[…]
51–75 -0.1268 Two or more 0.0291
[…]
76–100 -0.1286 Job Market Year (omitted)
(0.0665) [0.0710] ||0.0665|| Number of Letters (ref. One)
101–150 -0.0856 Two -0.2439
[…]
151–200 -0.1368∗∗ Three -0.1877
[…]
201+ -0.1473 Four -0.0817
[…]
Figure A.1: Prompt-based Sentiment Distribution for Different Letter Writers
Notes: NN=553 (complete applications sub-sample). The order of letter writers other than the main adviser has been randomized.
Table A.3: Predicting Success: complete applications sub-sample
[…]
Average Letter Length 0.2374∗∗∗ 0.2224∗∗∗ 0.2247∗∗∗ 0.2440∗∗∗ 0.1737∗∗∗
[…]
($p.$) (0.0077)∗∗∗ (0.0067)∗∗∗ (0.0079)∗∗∗ (0.0089)∗∗∗ (0.0075)∗∗∗
[0.0068]∗∗∗ [0.0064]∗∗∗ [0.0073]∗∗∗ [0.0092]∗∗∗ [0.0066]∗∗∗
[…]
Number of Letters Yes Yes Yes Yes Yes
Candidate’s characteristics Yes Yes Yes Yes Yes
PhD characteristics Yes Yes Yes Yes Yes
Job Market Year Yes Yes Yes Yes Yes
Number of top-5% Writers Yes Yes Yes Yes Yes
Adjusted $R^{2}$ 0.1194 0.1322 0.1210 0.1308 0.1008
[…]
Table A.4: Sentiment Dispersion and Job Market Success: alternative measures
Baseline Outcome
(1) (2) (3)
Average Letter Sentiment 0.0368∗∗∗ 0.0366∗∗∗ 0.0360∗∗∗
($p.$) (0.0087) (0.0087) (0.0087)
[0.0080] [0.0080] [0.0079]
||0.0107|| ||0.0108|| ||0.0108||
Sentiment Range (max-min) -0.0219∗∗∗
($p.$) (0.0072)
[0.0077]
||0.0073||
Sentiment Standard Deviation -0.0419∗∗∗
($p.$) (0.0143)
[0.0151]
||0.0147||
Sentiment Mean Absolute Deviation -0.0532∗∗∗
($p.$) (0.0191)
[0.0201]
||0.0200||
Four Letters & Avg. Letter Length Yes Yes Yes
Candidate’s characteristics Yes Yes Yes
PhD characteristics Yes Yes Yes
Job Market Year Yes Yes Yes
Number of top-5% Writers Yes Yes Yes
Adjusted $R^{2}$ 0.1324 0.1312 0.1296
[…]
@@ 3.1 Bert Score 4/16E 3
Max Depth 6
Min Elements to Split Node 21
Min Elements in Leaf Node 8
Forest Size 120
Notes: Hyper-parameters result from 5-fold CV search. Samples are bootstrapped and $\sqrt{13}\approx 4$ predictors are selected at random in the CART tree estimation. The impurity measure is the Gini. Observations are weighted by the inverse probability of the outcome to address class imbalance.
Figure A.2: The Distribution of Alternative Sentiment Measures
Note: NN=1968. Fitted line is a locally weighted regression.
Figure A.3: The Distribution of the Net-Standout Score
Note: NN=1968. Fitted line is a locally weighted regression.
Table A.6: Gendered Sentiment and Job Market Outcomes
[…]
Average RL Length 0.2141∗∗∗ 0.1770∗∗∗ 0.1754∗∗∗ 0.2109∗∗∗ 0.1578∗∗∗
(000s words) (0.0527) (0.0500) (0.0536) (0.0548) (0.0521)
[0.0563] [0.0463] [0.0577] [0.0580] [0.0575]
||0.0554|| ||0.0472|| ||0.0510|| ||0.0489|| ||0.0519||
[…]
[0.0080]∗∗ [0.0068] [0.0086]∗∗∗ [0.0091]∗∗∗ [0.0078]∗∗∗
[…]
Number of Letters & Avg. RL Length Yes Yes Yes Yes Yes
Candidate’s characteristics Yes Yes Yes Yes Yes
PhD characteristics Yes Yes Yes Yes Yes
Job Market Year Yes Yes Yes Yes Yes
Number of top-5% Writers Yes Yes Yes Yes Yes
Adjusted $R^{2}$ 0.1323 0.1322 0.1236 0.1467 0.1166
[…]